# Microscopic Model for Fractional Quantum Hall Nematics

### v2

@@ Abstract
[…]
${}^{1}$chool of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, United Kingdom
${}^{2}$epartment of Physics and Astronomy, The University of Tennessee, Knoxville, TN 37996, USA
${}^{3}$nstitute of Mathematical Sciences, CIT Campus, Chennai 600113, India
@@
${}^{4}$omi Bhabha National Institute, Training School Complex, Anushaktinagar, Mumbai 400094, India
@@
${}^{5}$epartment of Physics, University of Illinois at Urbana-Champaign, 1110 West Green Street, Urbana, Illinois 61801, USA
@@
${}^{6}$nthony J. Leggett Institute for Condensed Matter Theory, University of Illinois at Urbana-Champaign, 1110 West Green Street, Urbana, Illinois 61801, USA
@@
In this Supplemental Material, we present further details on the methodology used in the main text and additional results that support our conclusions: (i) we introduce three classes of variational wave functions that we have used to demonstrate the existence of FQH nematic transition in system sizes beyond exact diagonalization; (ii) we show that the FQH nematic transition can be probed by studying the dynamical response of the system to the geometric quench; (iii) we outline our procedure for calculating the Hall viscosity, which simplifies the method in Ref. [78]; (iv) we provide a detailed study of the charge-density-wave (CDW) phase in the torus geometry; (v) we demonstrate that our conclusions in the main text also hold for fermions at filling factor $\nu{=}1/3$.
SI Variational wave functions for the magnetoroton mode
@@
SI.1 Jack polynomial wave functions and Girvin-MacDonald-Platzman ansatz
@@
Here, $\mathrm{sl}_{\mu}$ denotes a Slater determinant with an orbital occupation pattern given by the partition $\mu$. The first condition thus means that we are looking for states that are obtained by squeezing from one of the roots $\lambda$ in Eq. (SI.1), which introduces a partial ordering among the partitions, denoted by $\preceq$. The second condition states that we are only interested in rotationally invariant states. These two conditions uniquely fix the coefficients $a_{\mu}$ of the Laughlin ground state wave function. For the excited magnetoroton states, to find $a_{\mu}$ we require the third, energetic condition, involving the $\hat{V}_{1}$ pseudopotential operator and electron annihilation operators $\hat{c}_{m}$, which singles out the state with quasiparticles piled up at the north pole of the sphere. The resulting states are confirmed to be excellent approximations to the eigenstates of Coulomb interaction [80]. Moreover, the construction can be readily generalized to other types of states that have a Jack polynomial description and a parent Hamiltonian, e.g., the Read-Rezayi sequence [101].
@@
At $L{=}1$, the density operator in Eq. (S5) reduces to the angular momentum raising operator [102], hence it annihilates the ground state, which is a singlet ($L{=}0$). Thus, the $L{=}1$ SMA trial state identically vanishes, which is also the case for the Jack construction. Moreover, at $L{=}2,3$, the SMA wave functions only involve the elements of the basis squeezed from the same root partition that defines the Jack model wave functions in Eq. (SI.1). Hence, for the Laughlin $1/3$ state, the equivalence between the two sets of wave functions can be proven for $L{=}2,3$ [80]. Moreover, both of these states are also equivalent to the CF exciton states discussed below. For $L{>}3$ the SMA wave functions contain unsqueezed basis components with respect to the root partitions used in the Jack model wave functions, thus the final states are no longer equivalent. The accuracy of the SMA ansatz in Eq. (S5), however, is numerically found to deteriorate rapidly at larger values of $L$ [103], unlike the Jack wave functions which remain accurate over the entire range $2{\leq}L{\leq}N$.
SI.2 LLL-projected GMP wave function
@@
SI.3 Composite fermion exciton wave function
@@
SII Geometric quench dynamics
@@
SIII Computation of the Hall viscosity
@@
SIV Charge density wave
@@
SV Fermionic Laughlin state

### v3

@@ Abstract
Songyang Pu1,2, Ajit C. Balram3,4, Joseph Taylor1, Eduardo Fradkin5,6 and Zlatko Papić1
1School of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, United Kingdom
2Department of Physics and Astronomy, The University of Tennessee, Knoxville, TN 37996, USA
3Institute of Mathematical Sciences, CIT Campus, Chennai 600113, India
@@
4Homi Bhabha National Institute, Training School Complex, Anushaktinagar, Mumbai 400094, India
@@
5Department of Physics, University of Illinois at Urbana-Champaign, 1110 West Green Street, Urbana, Illinois 61801, USA
@@
6Anthony J. Leggett Institute for Condensed Matter Theory, University of Illinois at Urbana-Champaign, 1110 West Green Street, Urbana, Illinois 61801, USA
@@
In this Supplemental Material, we present further details on the methodology used in the main text and additional results that support our conclusions: (i) we discuss in detail the different types of gaps that are studied in the main text; (ii) we introduce three classes of variational wave functions that we have used to demonstrate the existence of FQH nematic transition in system sizes beyond exact diagonalization; (iii) we show that the FQH nematic transition can be probed by studying the dynamical response of the system to the geometric quench; (iv) we show our main result about the existence of FQHN critical point is insensitive to the boundary condition by calculating the spectrum in disk geometry; (v) we outline our procedure for calculating the Hall viscosity, which simplifies the method in Ref. [76]; (vi) we provide a detailed study of the charge-density-wave (CDW) phase in the torus geometry; (vii) we demonstrate that our conclusions in the main text also hold for fermions at filling factor $\nu{=}1/3$.
SI Definitions of various gaps
In the main text, we have computed three types of gaps to characterize the fractional quantum Hall nematic (FQHN) transition: (i) the neutral gap is defined as the difference in the lowest two energies at the ground state flux, i.e., $22Q{=}2N{-}$ 2 for the $\nu_{b}{=}1/2$ Laughlin state; (ii) the charge or transport gap, i.e., the energy to create an individual quasihole (occurs at flux $12Q{+}$ 1) and quasiparticle (occurs at flux $12Q{-}$ 1) on top of the ground state; (iii) the gap $\Delta^{L{=}N}$ – the energy difference between the ground state and the lowest-lying state with orbital angular momentum $L{=}N$ – at the ground state flux. These three types of gaps were computed for the accessible system sizes $N{=}6{-}12$ bosons and their extrapolation to the thermodynamic limit in $1/N$ was performed. For the charge gap, it is important to include the neutralizing background charge correction [108, 61] Since we are considering modified Coulomb interaction, the total electrostatic energy is evaluated according to the general formula in the Supplementary Material of Ref. [109], which we reproduce here for convenience:
$E_{\mathrm{corr}}(2Q)=-\frac{1}{2}\mathcal{C}_{2Q}\left(N^{2}-n_{q}^{2}e_{*}^{2}\right),$ (S1)
$\mathcal{C}_{2Q}=\sum_{m=0}^{2Q}V_{m}\frac{4Q-2m+1}{(2Q+1)^{2}},$ (S2)
where $n_{q}$ is an integer taking values 0 (for the ground state) or ±${\pm}1$ for the quasiparticle and quasihole excitations that carry charge $e_{*}{=}e/2$ in our case. The energy correction explicitly depends on the number of flux quanta $2Q$ (we have assumed we are in the lowest Landau level). The average charging energy of the full LL per pair, $\mathcal{C}$, is determined by all pseudopotentials $V_{m}$, which includes any modification to the $V_{0}$. Furthermore, to minimize the finite-size effects, we multiply all energies with the density-correction factor $\sqrt{2Q\nu_{b}/N}$ [110]. In the case of charge gap, we do not density-correct the individual quasihole and quasiparticle energies but only apply the rescaling at the end.
We use the gap $\Delta^{L{=}N}$ as a cross-check for the extrapolated value of the charge gap in the thermodynamic limit. The $L{=}N$ magnetoroton state is formed by a quasiparticle and a quasihole on opposite poles of the sphere. As the charge gap is expected to remain finite across the FQHN transition, the quasiparticle and quasihole are localized objects with a few $\ell_{B}$ in diameter. By contrast, their interaction scales inversely with the diameter of the sphere, $2R{\propto}\sqrt{N}\ell_{B}$. Thus, in the thermodynamic limit, the quasiparticle and quasihole are infinitely far from each other and thus the interaction between them vanishes in this limit and they can be regarded as independent entities, the sum of whose energies is equivalent to the charge gap.
SII Variational wave functions for the magnetoroton mode
@@
SII.1 Jack polynomial wave functions and Girvin-MacDonald-Platzman ansatz
@@
Here, $\mathrm{sl}_{\mu}$ denotes a Slater determinant with an orbital occupation pattern given by the partition $\mu$. The first condition thus means that we are looking for states that are obtained by squeezing from one of the roots $\lambda$ in Eq. (SII.1), which introduces a partial ordering among the partitions, denoted by $\preceq$. The second condition states that we are only interested in rotationally invariant states. These two conditions uniquely fix the coefficients $a_{\mu}$ of the Laughlin ground state wave function. For the excited magnetoroton states, to find $a_{\mu}$ we require the third, energetic condition, involving the $\hat{V}_{1}$ pseudopotential operator and electron annihilation operators $\hat{c}_{m}$, which singles out the state with quasiparticles piled up at the north pole of the sphere. The resulting states are confirmed to be excellent approximations to the eigenstates of Coulomb interaction [95]. Moreover, the construction can be readily generalized to other types of states that have a Jack polynomial description and a parent Hamiltonian, e.g., the Read-Rezayi sequence [113].
@@
At $L{=}1$, the density operator in Eq. (S7) reduces to the angular momentum raising operator [114], hence it annihilates the ground state, which is a singlet ($L{=}0$). Thus, the $L{=}1$ SMA trial state identically vanishes, which is also the case for the Jack construction. Moreover, at $L{=}2,3$, the SMA wave functions only involve the elements of the basis squeezed from the same root partition that defines the Jack model wave functions in Eq. (SII.1). Hence, for the Laughlin $1/3$ state, the equivalence between the two sets of wave functions can be proven for $L{=}2,3$ [95]. Moreover, both of these states are also equivalent to the CF exciton states discussed below. For $L{>}3$ the SMA wave functions contain unsqueezed basis components with respect to the root partitions used in the Jack model wave functions, thus the final states are no longer equivalent. The accuracy of the SMA ansatz in Eq. (S7), however, is numerically found to deteriorate rapidly at larger values of $L$ [115], unlike the Jack wave functions which remain accurate over the entire range $2{\leq}L{\leq}N$.
SII.2 LLL-projected GMP wave function
@@
SII.3 Composite fermion exciton wave function
@@
SIII Geometric quench dynamics
@@
SIV Disk geometry
In the main text, we have focused on closed geometries, sphere and torus, which are most convenient for studying the system’s bulk properties and their extrapolation to the thermodynamic limit. In particular, we showed that our main conclusions about the existence of the FQHN critical point are largely insensitive to the choice of boundary condition. Hence, for sufficiently large systems, we expect these conclusions to hold also on a finite disk, which describes more closely the real physical system that necessarily has a boundary.
In Fig. S3 we confirm that the FQHN critical point is also realized on a finite disk in a similar parameter regime $\delta V_{0}{=}{-}0.4$. The FQH Hamiltonian on a disk conserves the total angular momentum $L_{z}$ about the perpendicular $z$-axis and the bosonic Laughlin state at $\nu_{b}{=}1/2$ occurs in the sector $L_{z}^{0}{=}N(N{-}1)$. However, block-diagonalization of the Hamiltonian in $L_{z}$-invariant sectors is insufficient to reveal the neutral spectrum [120]; instead, it is necessary to consider the approximately conserved center-of-mass angular momentum $\hat{M}_{\mathrm{COM}}$:
$\hat{M}_{\mathrm{COM}}=\hat{B}^{\dagger}\hat{B},\quad\hat{B}=\frac{1}{N}\sum_{i}\hat{R}_{i},$ (S24)
where $\hat{R}_{i}$ is the guiding center coordinate of the $i$th particle. This operator can be expressed in second quantization with the help of Clebsch-Gordan coefficients on the disk, which can be found, e.g., in Ref. [121]. For sufficiently small values of $L_{z}$, the spectrum of $\hat{M}_{\mathrm{COM}}$ is quantized in terms of integers [120]. The collective mode of a FQH state can be resolved by projecting the Hamiltonian into a subspace of states with eigenvalue 0 of $\hat{M}_{\mathrm{COM}}$ and then plotting the energies as a function of $\Delta L{\equiv}L_{z}^{0}{-}L_{z}$. To make the numerics more tractable, as customary in the literature, we also assume a cutoff $12N{-}$ 1 for the total number of orbitals accessible to the particles. We note that, due to the additional step of finding the kernel of $\hat{M}_{\mathrm{COM}}$, the diagonalization scheme on the disk geometry is significantly less efficient compared to the sphere or torus.
The presence of an edge can severely impact the bulk properties of small systems. Nevertheless, Fig. S3 shows that this is not the case here, and we can resolve the collective mode despite a fairly small number of particles. The softening of the Coulomb interaction by reducing $V_{0}$ has a similar effect as on the sphere and torus geometries: it makes the magnetoroton mode gapless in the long-wavelength limit, resulting in a nematic instability. In principle, by studying the momentum sectors $L_{z}{>}L_{z}^{0}$, the effect of nematic fluctuations on FQH edge states could also be explored in this setup, which we leave for future work.
Figure S3: Energy spectrum for bosons at $\nu_{b}{=}1/2$ on the disk for pure Coulomb interaction (red circles) and softened Coulomb interaction with $\delta V_{0}{=}{-}0.4$ (blue diamonds). Data is for $N{=}7$ bosons in 13 orbitals with the center-of-mass momentum equal to zero. The spectrum is plotted relative to the momentum of the Laughlin ground state (see text). Similar to the sphere and torus, the softening of $V_{0}$ pseudopotential makes the magnetoroton mode (dashed line) go soft in the long-wavelength limit (dotted line). The lines are a guide to the eye.
@@
SV Computation of the Hall viscosity
@@
SVI Charge density wave
@@
SVII Fermionic Laughlin state